(t-4)^2 -32 = (t+4)^2
Answer by htmentor(1343)

I assume we are supposed to solve for t?
When we apply FOIL to the quadratic terms, the t^2 and the constant term, 16, will cancel:
(t-4)^2 -32 = (t+4)^2 -> -8t - 32 = 8t
16t = -32
This gives t = -2
Check:
(-6)^2 - 32 = 4 = (2)^2